The tangent to the curve 𝑦 = 2π‘₯2 βˆ’ 5π‘₯ + 6 at the point (2,1) intersects the normal to the same curve at the point (1,4) at point 𝑄. Find the coordinates of 𝑄


1
Expert's answer
2022-03-03T18:06:55-0500
"y=2x^2-5x+6"

Point "(2,1)" does not lie on the curve. Point "(1,4)" does not lie on the curve.


The tangent to the curve "y=2x^2-5x+6" at the point "(2,4)" intersects the normal to the same curve at the point "(1,3)" at point "\ud835\udc44." Find the coordinates of "Q."


"y'=4x-5"

Point "(2,4)"


"slope_1=m_1=4(2)-5=3"

The equation of the tangent line at "(2, 4)"


"y=3(x-2)+4"

"y=3x-2"

Point "(1,3)"

"slope_2=m_2=-\dfrac{1}{3(1)-5}=-\dfrac{1}{2}"

The equation of the normal line at "(1, 4)"

"y=-\dfrac{1}{2}(x-1)+3"

"y=-\dfrac{1}{2}x+\dfrac{7}{2}"


Then


"3x-2=-\dfrac{1}{2}x+\dfrac{7}{2}"

"x=\dfrac{11}{7}"

"y=3(\dfrac{11}{7})-2=\dfrac{19}{7}"

"Q\ (\dfrac{11}{7}, \dfrac{19}{7})"